A monopolist can produce any level of output at a constant marginal cost of $5 per unit. Assume the monopoly sells its goods in two different markets separated by some distance. The demand curve in the first market is given by q1 = 65 − p1,and the demand curve in the second market is given by q2 = 90 − 2p2.

  1. (a) If the monopolist can maintain the separation between the two markets, what level of output should be produced in each market, and what price will prevail in each market? What are the total profits in this situation?

  2. (b) What is the price elasticity of demand at the monopoly output in each market?

  3. (c) Now suppose the monopolist cannot maintain separation between the two mar- kets. Specifically, it costs consumers only $4 to transport goods between the two markets. What would be the monopolist’s new profit level in this situation?

  4. (d) How would your answer to part (c) change if the transportation costs were zero?
Answer #1

we have the following information - Mkt 1 dd: 9 = 65-P1 ZA MC = MC2 = 5 Mkt 2 dd; 92= 90-282} MC;= des a a. e, = ssdgia sorot

To market i avi = 68-pill 0912-1 opo lejl= dr x ha-1 x 35 til 117 i dpi ali on 30 For mobet 2; to 922 90-292 day2=-2 op2 Tezl